# Implicit Neural Representations for Robust Joint Sparse-View CT Reconstruction

### v1

@@ 4.2 Algorithm Implementation
where $\widehat{\bm{w}}_{j}$ denotes a sample from $q(\bm{w}_{j})$. We only do MC sampling once at each iteration, which works efficiently in practice. To facilitate the gradient-based optimization schemes, we utilize the reparameterization trick (Kingma & Welling, 2013): $q(\bm{w}_{j})=\bm{\mu}_{j}+\log\left(1+\exp{(\bm{\pi}_{j})}\right)\mathcal{N}(\bm{0},\bm{I})$. Here, we additionally employ the softplus function in parameterizing the variance $\bm{\sigma}_{j}$ with the variable $\bm{\pi}_{j}$ to ensure the non-negativity of the variance.
@@ 5 Experiments
Comparison Methods. We compare our approach with other methods that do not require a set of densely measured data, thereby excluding those of supervised learning: i) Classical techniques: Filtered Back Projection (FBP) and Simultaneous Iterative Reconstruction Technique (SIRT); ii) Naive INR-based single reconstruction method, denoted as SingleINR; iii) FedAvg, a federated averaging approach proposed by Kundu et al. (2022); iv) MAML, a meta-learning technique as discussed by Tancik et al. (2021); v) INRWild, a method adapted from NeRFWild (Martin-Brualla et al., 2021). FBP and SIRT are classical methods that do not use networks, while all other methods employ an identical INR network as described in the next paragraph.
INR Network Configuration. The same backbone and associated configurations are applied to ensure a fair comparison. All INR-based methods employ the SIREN architecture (Sitzmann et al., 2020) coupled with the same positional embedding (Tancik et al., 2020). Specifically, we consistently use a Siren network with the following dimensions: a depth of 8 and width of 128 for the WalnutCT and AluminiumCT experiments, and a depth of 8 with width 256 for the rest experiments. Position encoding is implemented using Fourier feature embedding (Tancik et al., 2020), with an embedding dimension of 256 for WalnutCT and AluminiumCT, 512 for the rest. Consistently, this embedded position is input to the INRs to facilitate density prediction. In alignment with Martin-Brualla et al. (2021), INRWild utilizes an 8-layer SIREN network for the static segment and a 4-layer SIREN for each transient component.
@@ 5.1 Results
Experiment Metrics FBP SIRT SingleINR INRWild MAML FedAvg INR-Bayes
Inter-object PSNR $19.54\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.25}}$ $23.82\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.25}}$ $35.18\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.43}}$ $19.76\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.35}}$ $35.66\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.38}}$ $31.57\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.32}}$ $\textbf{36.13}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.33}}$
Walnut SSIM $0.328\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.003}}$ $0.435\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.007}}$ $0.934\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.004}}$ $0.675\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.010}}$ $0.941\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.003}}$ $0.886\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.004}}$ $\textbf{0.946}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.003}}$
4DCT PSNR $25.19\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.03}}$ $28.61\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.03}}$ $34.07\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.04}}$ $33.76\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.04}}$ $34.35\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.05}}$ $34.67\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.04}}$ $\textbf{34.79}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.04}}$
SSIM $0.529\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.746\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.877\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.863\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.881\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.885\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $\textbf{0.894}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$
Intra-object PSNR $26.50\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$ $28.81\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$ $32.80\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.11}}$ $28.46\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.07}}$ $33.26\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.10}}$ $32.42\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.08}}$ $\textbf{33.90}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.10}}$
Lung SSIM $0.568\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.719\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.815\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.674\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.003}}$ $0.825\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.808\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $\textbf{0.840}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$
Inter-object PSNR $32.13\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.05}}$ $28.61\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.33}}$ $35.62\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$ $29.77\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.05}}$ $36.22\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$ $35.75\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$ $\textbf{36.67}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$
Aluminium SSIM $0.717\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.836\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.850\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.694\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.871\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.849\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $\textbf{0.881}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$
Inter-object PSNR $24.97\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.14}}$ $28.32\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.13}}$ $32.64\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.27}}$ $25.05\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.18}}$ $33.13\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.22}}$ $31.68\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.19}}$ $\textbf{33.75}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.20}}$
Lung SSIM $0.503\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.004}}$ $0.678\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.005}}$ $0.821\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.007}}$ $0.560\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.08}}$ $0.833\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $0.807\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $\textbf{0.847}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$
Inter-object PSNR $18.12\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.39}}$ $29.13\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.20}}$ $30.90\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.32}}$ $16.43\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.28}}$ $25.71\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.27}}$ $30.74\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.29}}$ $\textbf{31.31}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.31}}$
Faces SSIM $0.549\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $0.766\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.005}}$ $0.831\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.008}}$ $0.260\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.011}}$ $0.647\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.012}}$ $0.827\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.007}}$ $\textbf{0.847}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.007}}$
@@
Table 1: Performance comparison across different settings and datasets. The highest average PSNR/SSIM values are highlighted in bold, consistently showing that INR-Bayes outperforms other methods. Reconstruction of human faces is included to illustrate our method’s broad applicability, despite lacking practical relevance in physical contexts.
@@
Reconstruction Performance. Table 1 shows that our method consistently achieves the highest average metrics across datasets, excelling in leveraging inherent trends in slices. This advantage is particularly pronounced in experiments with distinct transition patterns, as seen in both inter-object and intra-object settings. In contrast, INRWild generally underperforms compared to SingleINR. Its strategy of jointly reconstructing static representations while independently capturing transient characteristics fails to enhance overall reconstruction quality. In most scenarios, FedAvg is outperformed by SingleINR, suggesting its inability to learn a meta representation that meaningfully improves reconstruction quality. MAML ranks as the second-best method in most settings, indicating its learned meta representation aids in individual reconstruction performance.
@@
The visual comparisons in Figure 3 further validate our results. Reconstructions by SingleINR exhibit noticeable artifacts, while FedAvg and MAML struggle to capture finer image details. In contrast, INR-Bayes achieves superior visual quality, effectively balancing smoothness and detail preservation, yielding results closer to the ground truth.
@@
Figure 3: Visual comparison of reconstruction performance on all used datasets. Enlarged areas are highlighted in red insets. PSNR values are on the top left, with SSIM values on the bottom left. Reconstruction of human faces is included to illustrate our method’s broad applicability, despite lacking practical relevance in physical contexts. Illustrations figures (first column) are modified from J.Dncsn (2009); Agashi5859 (2014); Lene (2021); Liu et al. (2015).
@@
Overfitting. Iterative reconstruction methods tend to overfit when applied to limited data (Herman & Odhner, 1991; Effland et al., 2020). In contrast, Bayesian frameworks have demonstrated robustness against overfitting (MacKay, 1992; Neal, 2012; Blundell et al., 2015). To validate this, we extend the training iterations in Table 1 from 30K to 60K. As shown in Figure 4 on inter-object LungCT, the learning curves of baselines deteriorate in the long run, indicating overfitting during optimization. Conversely, our approach maintains a consistent level of reconstruction quality once the optimal performance is achieved, underscoring its robustness.
@@
Challenge in Determinating Appropriate Stopping Criterion. We find that in the noisy CT reconstruction environments it is inherently challenging to determine the optimal stopping point when the overfitting problem presents. This is due to the variation of convergence speed across different objects. To demonstrate that, we present the individual training curves from three different patients within the noisy LungCT dataset in Figure 7. Despite originating from the same dataset, these three cases exhibit distinctly different optimal stopping points, suggesting that a one-size-fits-all approach to training termination is ineffective for appraches suffering from the overfitting problem.
Figure 7: Training curves for three different patients from the noisy LungCT dataset, each showing unique optimal stopping points. This variability exemplifies the challenges in universally determining the precise moment to stop training.

### v2

@@ 4.2 Algorithm Implementation
where $\widehat{\bm{w}}_{j}$ denotes a sample from $q(\bm{w}_{j})$. We only do MC sampling once at each iteration, which works efficiently in practice. To facilitate the gradient-based optimization schemes, we utilize the reparameterization trick (Kingma & Welling, 2013): $q(\bm{w}_{j})=\bm{\mu}_{j}+\log\left(1+\exp{(\bm{\pi}_{j})}\right)\mathcal{N}(\bm{0},\bm{I})$. Here, we additionally employ the softplus function in parameterizing the variance $\rho_{j}$ with the variable $\bm{\pi}_{j}$ to ensure the non-negativity of the variance.
@@ 5 Experiments
Comparison Methods. We compare our approach with other methods that do not require a set of densely measured data, thereby excluding those of supervised learning: i) Classical techniques: Filtered Back Projection (FBP) and Simultaneous Iterative Reconstruction Technique (SIRT) (Gilbert, 1972); ii) Regularization-based approaches with model-based reconstruction method FISTA (Beck & Teboulle, 2009): total generalized variation denoted as RegTGV proposed by Niu et al. (2014), wavelet based regularization RegWavelet (Zhang et al., 2018a; Garduño et al., 2011) and second order Lysaker-Lundervold-Tai with Rudin-Osher-Fatemi TV, denoted as RegLLT-TV (Kazantsev et al., 2017); iii) Naive INR-based single reconstruction method, denoted as SingleINR; iv) FedAvg, a federated averaging approach proposed by Kundu et al. (2022); v) MAML, a meta-learning technique as discussed by Tancik et al. (2021); vi) INRWild, a method adapted from NeRFWild (Martin-Brualla et al., 2021). FBP and SIRT are classical methods that do not use networks, while all other methods employ an identical INR network as described in the next paragraph.
INR Network Configuration. The same backbone and associated configurations are applied to ensure a fair comparison. All INR-based methods employ the SIREN architecture (Sitzmann et al., 2020) coupled with the same positional embedding (Tancik et al., 2020). Specifically, we consistently use a Siren network with the following dimensions: a depth of 8 and width of 128 for the WalnutCT and AluminiumCT experiments, and a depth of 8 with width 256 for the rest of the experiments. Position encoding is implemented using Fourier feature embedding (Tancik et al., 2020), with an embedding dimension of 256 for WalnutCT and AluminiumCT, 512 for the rest. Consistently, this embedded position is input to the INRs to facilitate density prediction. In alignment with Martin-Brualla et al. (2021), INRWild utilizes an 8-layer SIREN network for the static segment and a 4-layer SIREN for each transient component. We note that all nodes use the same size network for our method and comparison methods.
@@ 5.1 Results
Method Metrics Inter-walnut 4DCT Intra-lung Inter-Aluminium Inter-Lung Inter-Faces
FBP PSNR $19.54\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.25}}$ $25.19\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.03}}$ $26.50\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$ $32.13\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.05}}$ $24.97\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.14}}$ $18.12\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.39}}$
SSIM $0.328\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.003}}$ $0.529\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.568\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.717\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.503\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.004}}$ $0.549\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$
SIRT PSNR $23.82\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.25}}$ $28.61\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.03}}$ $28.81\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$ $28.61\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.33}}$ $28.32\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.13}}$ $29.13\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.20}}$
SSIM $0.435\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.007}}$ $0.746\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.719\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.836\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.678\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.005}}$ $0.766\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.005}}$
RegTGV PSNR $30.64\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.33}}$ $31.50\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.04}}$ $31.46\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.07}}$ $27.21\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.03}}$ $31.18\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.19}}$ $28.39\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.18}}$
SSIM $0.872\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.005}}$ $0.859\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.810\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.868\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.809\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $0.833\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.005}}$
RegWavelet PSNR $24.79\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.25}}$ $27.00\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.03}}$ $27.95\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.05}}$ $25.75\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.03}}$ $27.76\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.11}}$ $27.00\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.21}}$
SSIM $0.460\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $0.594\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.634\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.817\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.626\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.004}}$ $0.701\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$
RegLLT-TV PSNR $33.33\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.30}}$ $32.39\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.04}}$ $32.26\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.08}}$ $27.19\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.03}}$ $31.95\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.19}}$ $30.09\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.22}}$
SSIM $0.933\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.890\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.828\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.867\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.828\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $\textbf{0.858}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.005}}$
SingleINR PSNR $35.18\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.43}}$ $34.07\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.04}}$ $32.80\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.11}}$ $35.62\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$ $32.64\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.27}}$ $30.90\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.32}}$
SSIM $0.934\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.004}}$ $0.877\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.815\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.850\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.821\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.007}}$ $0.831\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.008}}$
INRWild PSNR $19.76\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.35}}$ $33.76\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.04}}$ $28.46\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.07}}$ $29.77\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.05}}$ $25.05\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.18}}$ $16.43\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.28}}$
SSIM $0.675\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.010}}$ $0.863\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.674\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.003}}$ $0.694\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.560\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.08}}$ $0.260\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.011}}$
MAML PSNR $35.66\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.38}}$ $34.35\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.05}}$ $33.26\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.10}}$ $36.22\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$ $33.13\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.22}}$ $25.71\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.27}}$
SSIM $0.941\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.003}}$ $0.881\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.825\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.871\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.833\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $0.647\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.012}}$
FedAvg PSNR $31.57\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.32}}$ $34.67\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.04}}$ $32.42\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.08}}$ $35.75\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$ $31.68\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.19}}$ $30.74\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.29}}$
SSIM $0.886\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.004}}$ $0.885\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.808\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $0.849\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $0.807\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $0.827\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.007}}$
INR-Bayes PSNR $\textbf{36.13}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.33}}$ $\textbf{34.79}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.04}}$ $\textbf{33.90}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.10}}$ $\textbf{36.67}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.06}}$ $\textbf{33.75}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.20}}$ $\textbf{31.31}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.31}}$
SSIM $\textbf{0.946}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.003}}$ $\textbf{0.894}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $\textbf{0.840}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.002}}$ $\textbf{0.881}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $\textbf{0.847}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $0.847\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.007}}$
@@
Table 1: Performance comparison across different settings on noiseless measurements. The highest average PSNR/SSIM values are highlighted in bold, consistently showing that INR-Bayes outperforms other methods. Reconstruction of human faces is included to illustrate our method’s broad applicability, despite lacking practical relevance in physical contexts.
@@
Reconstruction Performance. Table 1 shows that our method consistently achieves the highest average metrics across datasets on noiseless measurements, with the exception of the SSIM values on the CelebA dataset, excelling in leveraging inherent trends in slices. This advantage is particularly pronounced in experiments with distinct transition patterns, as seen in both inter-object and intra-object settings. In contrast, INRWild generally underperforms compared to SingleINR. Its strategy of jointly reconstructing static representations while independently capturing transient characteristics fails to enhance overall reconstruction quality. In most scenarios, FedAvg is outperformed by SingleINR, suggesting its inability to learn a meta representation that meaningfully improves reconstruction quality. MAML ranks as the second-best method in most settings, indicating its learned meta representation aids in individual reconstruction performance. However, regularization-based methods typically underperform compared to INR-based methods, except on the CelebA dataset where RegLLT-TV achieves the highest SSIM value. This underscores the efficacy of the heuristic LLT-TV prior for natural images as opposed to CT images.
@@
The visual comparisons in Figure 3 further validate our results. Reconstructions by SingleINR exhibit noticeable artifacts, while FedAvg and MAML struggle to capture finer image details. RegLLT-TV often results in overly smooth reconstructions, but in contrast, INR-Bayes achieves superior visual quality, effectively balancing smoothness and detail preservation, yielding results closer to the ground truth.
@@
Figure 3: Visual comparison of reconstruction performance on noiseless measurements. Enlarged areas are highlighted in red insets. PSNR values are on the top left, with SSIM values on the bottom left. Reconstruction of human faces is included to illustrate our method’s broad applicability, despite lacking practical relevance in physical contexts. Illustrations figures (first column) are modified from J.Dncsn (2009); Agashi5859 (2014); Lene (2021); Liu et al. (2015).
@@
Overfitting. [^3] In the context of CT reconstruction, conventional iterative methods like SIRT typically approach a good approximation to the exact solution early in iterations and subsequently diverge from it, often attributed to convergence issues (Elfving et al., 2014). In contrast, in the machine learning domain, overfitting refers to a scenario where a neural network excessively fits to training data, leading to a drop in actual performance despite decreasing training loss (Srivastava et al., 2014). Our experiments contain both non-learning-based methods (SIRT) and learning-based methods (INR). We acknowledge the subtle differences between convergence challenges and overfitting but use the term overfitting broadly to describe performance deterioration in both noiseless and noisy scenarios. Iterative reconstruction methods tend to overfit when applied to limited data (Herman & Odhner, 1991; Effland et al., 2020) (particularly noticeable in sparse-view CT scenarios). In contrast, Bayesian frameworks have demonstrated robustness against overfitting (MacKay, 1992; Neal, 2012; Blundell et al., 2015). To validate this, we extend the training iterations in Table 1 from 30K to 60K. As shown in Figure 4 on inter-object LungCT, the learning curves of baselines deteriorate in the long run, indicating overfitting during optimization. Conversely, our approach maintains a consistent level of reconstruction quality once the optimal performance is achieved, underscoring its robustness.
@@
Challenge in Determinating Appropriate Stopping Criterion. We find that in the noisy CT reconstruction environments it is inherently challenging to determine the optimal stopping point when the overfitting problem presents. This is due to the variation of convergence speed across different objects. To demonstrate that, we present the individual SingleINR training curves from three different patients within the noisy LungCT dataset in Figure 7. Despite originating from the same dataset, these three cases exhibit distinctly different optimal stopping points, suggesting that a one-size-fits-all approach to training termination is ineffective for appraches suffering from the overfitting problem.
Figure 7: SingleINR training curves for three different patients from the noisy LungCT dataset, each showing unique optimal stopping points. This variability exemplifies the challenges in universally determining the precise moment to stop training.
@@
RegTGV $27.76\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.31}}$ $0.748\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.010}}$ $27.44\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.17}}$ $0.619\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $27.14\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.03}}$ $0.862\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $27.97\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.18}}$ $0.800\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.004}}$
RegWavelet $23.49\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.25}}$ $0.393\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $26.52\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.15}}$ $0.548\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.005}}$ $25.19\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.03}}$ $0.716\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $22.83\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.26}}$ $0.471\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.010}}$
RegLLT-TV $28.51\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.31}}$ $0.817\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.010}}$ $30.58\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.17}}$ $0.783\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.006}}$ $27.23\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.03}}$ $0.863\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$ $28.00\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.18}}$ $\textbf{0.801}\stackrel{{\scriptstyle}}{{\scriptstyle\pm 0.001}}$